OpenStudy (anonymous):

Help.... A man has two more nickels than dimes. He has $1.15 in all. How many coins of each kind does he have? Which equation can be used to solve the problem?

OpenStudy (anonymous):

Let n = number of nickels d = number of dimes n = 2 + d (equation 1) .05n + .10d = 1.15 (equation 2) Substitute equation 1 into equation 2: .05(2 + d) + .10d = 1.15 .10 + .05d + .10d = 1.15 .10 + .15d = 1.15 .15d = 1.05 d = 7 n = 2 + d n = 2 + 7 n = 9 The man has 9 nickels and 7 dimes

OpenStudy (anonymous):

THANKX